# Meta-LoRA: Meta-Learning LoRA Components for Domain-Aware ID Personalization

### v2

@@ 12 Baseline Model Details and Implementation Settings
We compare our Meta-LoRA personalization approach against several state-of-the-art subject-driven generation methods with publicly available implementations: InstantID [30], PhotoMaker [17], and PuLID [9]. Although each method’s GitHub demos recommend different diffusion backbones, to ensure a fair evaluation we run all SDXL-based methods under three popular variants:
@@
In contrast, our default LoRA and Meta-LoRA approaches are built on a FLUX-based model.
@@
All SDXL-based experiments are conducted on a single NVIDIA RTX 3090 GPU with 24 GB VRAM, while the Flux-based PuLID inference (and our FLUX-based LoRA and Meta-LoRA) are performed on a single NVIDIA A6000 GPU with 48 GB VRAM (using CPU offload). In every case, we use bf16 as the data type for inference.
PuLID (SDXL and FLUX).
We adopt the newest PuLID Flux implementation (FLUX_PuLID_v0.9.1) from its public repository, which is claimed to improve upon the original paper’s version. For the SDXL-based PuLID model, we set ID_scale=0.8 following the repository’s recommended configuration. For the Flux version, we set id_weight=1.0, true_cfg=1.0, time_to_start_cfg=1, inference_steps=20, cfg=4.0, and max_sequence_length=128, in accordance with the demo code.
PhotoMaker.
We use PhotoMakerV2, which differs from the version evaluated in the PuLID paper [9] (where its performance was reported as poor, and claimed it is not even compatible with SDXL-Lightning). In our experiments, PhotoMakerV2 yields substantially improved results—possibly owing to internal updates. Although PhotoMakerV2 typically expects SDXL-RealVis4.0 as its default backbone, we also evaluate it with SDXL-Lightning and SDXL-Base. To ensure compatibility with SDXL-Lightning (particularly for a 4-step inference), we set time_spacing="trailing" in the Euler scheduler configuration and adjust the start_merge_step parameter to 0 for SDXL-Lightning or 10 for SDXL-Base and SDXL-RealVis4.0. We further prepend the keyword “img” to each prompt (e.g., “man img ...”, “woman img ...”) to satisfy PhotoMaker’s trigger-word requirement.
InstantID.
For InstantID, we consider two implementations:
  1. 1.
    The ComfyUI-based approach used in the PuLID paper [9], which exposes a “denoise” parameter in the kSampler node (with values between 0 and 1).
  2. 2.
    The original InstantID GitHub repository’s implementation.
@@
Prompting and Negative Prompts for SDXL Models.
All SDXL-based methods (InstantID, PhotoMaker, and PuLID-SDXL) share the following negative prompt:
@@
“flaws in the eyes, flaws in the face, flaws, lowres, non-HDRi, low quality, worst quality, artifacts noise, text, watermark, glitch, deformed, mutated, ugly, disfigured, hands, low resolution, partially rendered objects, deformed or partially rendered eyes, deformed, deformed eyeballs, cross-eyed, blurry”
Flux-based methods (including PuLID-Flux, default LoRA, and Meta-LoRA) do not use any negative prompts, consistent with their default configurations.
SDXL Inference Details.
We employ the Euler scheduler for all SDXL inferences:
@@
  • SDXL-Lightning. We adopt the 4-step sdxl-lightning checkpoint from ByteDance’s HuggingFace page, with CFG-scale=1.2, following recommendations from the PuLID paper [9].
  • SDXL-Base 1.0 and SDXL-RealVis 4.0. We set CFG-scale=5.0 and use 50 inference steps, in line with standard SDXL configurations.
LoRA Baseline.
@@
We also include a standard LoRA fine-tuning baseline [10] (implemented on a FLUX-based model) as a reference. A rank-1 LoRA is trained for each identity. Initially, we train for 375 iterations (to match Meta-LoRA’s Stage-2 steps), but observe that identity fidelity remains underfitted. We therefore extend training to 675 iterations and apply data augmentation strategies. Although these extended and augmented LoRAs show improved performance, our Meta-LoRA—with its meta-learned prior—converges faster and achieves higher identity fidelity in fewer steps.
[…]
InstantID [30] SD-XL [21] 80.17 71.33 -11.0%
SD-XL Lightning [18] 82.91 74.77 -9.8%
RealVisXL V4 [26] 81.11 73.10 -9.9%
PhotoMaker [17] SD-XL 67.28 67.16 -0.2%
SD-XL Lightning 66.67 63.02 -5.5%
RealVisXL V4 68.94 68.57 -0.5%
PuLID [9] SD-XL 78.84 71.06 -9.7%
SD-XL Lightning 79.90 73.50 -8.0%
RealVisXL V4 78.78 75.19 -4.6%
FLUX.1-dev [15] 84.76 75.72 -10.7%
Rank-1 LoRA 375 Iters [10] FLUX.1-dev 60.60 57.62 -4.9%
+ augmented inputs FLUX.1-dev 70.55 72.69 +3.0%
Rank-1 LoRA 625 Iters FLUX.1-dev 64.32 61.41 -4.2%
+ augmented inputs FLUX.1-dev 76.77 76.15 -0.8%
Meta-LoRA FLUX.1-dev 79.45 77.16 -2.9%
Table 2: Comparison of Robust Face Sim and Face Sim metrics, with percentage difference highlighted for significant drops.
Input:
$I$ Set of identities with multiple images
$P$ Set of carefully chosen prompts
$M$ Generative text-to-image model
$C$ Face cropper
$E$ Identity embedding extractor
Output: $R\text{-}FaceSim$: Robust Face Similarity Score
for each $i\in I$ do
       $ref\leftarrow$ select_one($i$);
       $T\leftarrow i\setminus\{ref\}$;
      $M\leftarrow$ fine-tune/image-prompt($M$, $ref$);
      for each $p\in P$ do
             $g\leftarrow M(p)$;
             $f_{g}\leftarrow C(g)$;
             $e_{g}\leftarrow E(f_{g})$;
            $sumsim\leftarrow 0$, $count\leftarrow 0$;
             for each $t\in T$ do
                   $f_{t}\leftarrow C(t)$;
                   $e_{t}\leftarrow E(f_{t})$;
                   $sim\leftarrow cosine(e_{g},e_{t})$;
                   $sumsim\leftarrow sumsim+sim$;
                   $count\leftarrow count+1$;
            $S(i,p)\leftarrow\frac{sumsim}{count}$;
$R\text{-}FaceSim\leftarrow mean\{S(i,p):i\in I,p\in P\}$;
return $R\text{-}FaceSim$
Algorithm 3 Robust Face Similarity Metric (R-FaceSim)
@@
R-FaceSim. The algorithm 3 computes the Robust Face Similarity Metric (R-FaceSim) by leveraging multiple images per identity, carefully chosen prompts, and a generative text-to-image model. For each identity, a single reference image is selected while the remaining images form the test set. The model is conditioned or fine-tuned using the reference image. For each prompt, the model generates an output image from which the face is cropped and its identity embedding extracted. This embedding is then compared via cosine similarity to the embeddings of faces from each test image. The mean similarity for a given prompt and identity is computed by averaging these cosine similarities. Finally, R-FaceSim is obtained by averaging the prompt-level similarities across all identities and prompts.
@@
12.1 Evaluation Prompts
Table 3 and 4 include the prompts used in metric evaluations, used only with the Meta-PHD-FFHQ dataset. All metrics except R-FaceSim were calculated on these prompts.
The R-FaceSim metrics were calculated only on the Meta-PHD-UNSPLASH. Used for no other calculations. Designed specifically to provoke models to produce images where the person is facing the camera and no objects occluding their face.

### v3

@@ Appendix C Baseline Model Details and Implementation Settings
We evaluate our Meta-LoRA personalization approach against several state-of-the-art subject-driven generation methods with publicly available implementations: InstantID [29], PhotoMaker [17], and PuLID [9]. As a FLUX.1-dev-based [15] implementation of PuLID is available, which is claimed to improve upon the original paper’s version (i.e., FLUX_PuLID_v0.9.1 from its public repository [^2] https://github.com/ToTheBeginning/PuLID), we use this version for a direct comparison. For InstantID and PhotoMaker, although their respective GitHub demos recommend different diffusion backbones, we standardize the evaluation by running all SDXL-based methods under these two base variants to ensure fairness:
@@
In contrast, our default LoRA and Meta-LoRA approaches are built on a FLUX.1-dev-based [15] model.
@@
All experiments based on SDXL are conducted on a single NVIDIA RTX 3090 GPU with 24GB VRAM. In contrast, experiments involving FLUX.1-dev models—including PuLID, LoRA, and Meta-LoRA—are run on a single NVIDIA A6000 GPU with 48GB VRAM, utilizing CPU offloading. Additionally, LoRA training and the stage-2 training of Meta-LoRA (used solely for rank-based performance analysis) are carried out on a single NVIDIA H100 GPU with 66GB VRAM. In all cases, bf16 is used as the inference data type.
Model-specific inference configurations.
For Meta-LoRA and each of the studies that we provide a comparison, the following configurations are selected for inference:
  • Meta-LoRA and LoRA: inference_steps=30, cfg=3.5, true_cfg=1.0, lora_strength=1.0
  • PuLID: We follow the repository’s demo code, and set id_weight=1.0, true_cfg=1.0, time_to_start_cfg=1, inference_steps=20, cfg=4.0, and max_sequence_length=128.
  • PhotoMaker: We use PhotoMakerV2, which differs from the version evaluated in the PuLID paper [9] (where its performance was reported as poor, and claimed it is not even compatible with SDXL-Lightning). In our experiments, PhotoMakerV2 yields substantially improved results—possibly owing to internal updates. To ensure compatibility with SDXL-Lightning (particularly for a 4-step inference), we set time_spacing="trailing" in the Euler scheduler configuration and adjust the start_merge_step parameter to 0 for SDXL-Lightning or 10 for SDXL-Base. We further prepend the keyword “img” to each prompt (e.g., “man img ...”, “woman img ...”) to satisfy PhotoMaker’s trigger-word requirement.
@@
Common inference configurations.
Below, the configuration parameters are shared, which are common to multiple models:
  • Image size: We set the generated image size for all experiments (except InstantID) as 1024$\times$1024.
@@
  • Negative prompts: While FLUX.1-dev-based models do not leverage any negative prompts, the following text is used for all SDXL-based methods: "flaws in the eyes, flaws in the face, flaws, lowres, non-HDRi, low quality, worst quality, artifacts noise, text, watermark, glitch, deformed, mutated, ugly, disfigured, hands, low resolution, partially rendered objects, deformed or partially rendered eyes, deformed, deformed eyeballs, cross-eyed, blurry".
@@
  • SDXL inference details: We employ the Euler scheduler for all SDXL inferences. For SDXL-Lightning, we adopt the 4-step sdxl-lightning checkpoint from ByteDance’s HuggingFace page, with CFG-scale=1.2, following recommendations from the PuLID paper [9]. SDXL-Base 1.0, we set CFG-scale=5.0 and use 50 inference steps, in line with standard SDXL configurations.
Appendix D Choices of Rank and Iteration Count on Meta-LoRA and LoRA Trainings
On the rank selection.
Although we present results across various rank values in Figure 3 (Left) of the main paper, all of our qualitative and additional quantitative analyses are conducted using rank-1 configurations. This decision is motivated by several considerations:
  • Our goal is to maintain a minimal and straightforward setup, while still demonstrating that the model can effectively learn identity representations. To align with this objective, we also restrict fine-tuning to a single identity image.
  • With only one image available during fine-tuning, larger model capacities (i.e., higher ranks) increase the risk of overfitting. Rank-1 offers a good trade-off between CLIP-T and R-FaceSim performance, making it a practical and effective default choice for our study. See Section I for a more detailed analysis.
That said, our framework naturally supports scenarios involving multiple identity images during fine-tuning, which can be further explored in future work. In such cases, using higher ranks may further improve and stabilize model performance.
On the number of iterations.
As discussed in Section A, we evaluate our models across different rank values and training iterations. For Meta-LoRA, we empirically observed that iteration 375 achieves a well-balanced trade-off between prompt adherence and identity preservation—that is, the generated images accurately reflect the target identity without exhibiting noticeable overfitting to the input image. Thanks to its Stage-1 preconditioning on identity-related information, Meta-LoRA converges more rapidly than the default LoRA architecture. In contrast, the default LoRA model reaches a similar balance around iteration 625 (i.e., identity fidelity remains underfitted at earlier iterations). Based on these findings, we select iteration 375 for Meta-LoRA Stage-2 and iteration 625 for LoRA in our main evaluations.
@@ Appendix D Choices of Rank and Iteration Count on Meta-LoRA and LoRA Trainings
Although these extended and augmented LoRAs show improved performance, our Meta-LoRA—with its meta-learned prior—converges faster and achieves higher identity fidelity in fewer steps.
@@
Appendix E Details on Evaluation Datasets and Prompts
The FFHQ partition of the Meta-PHD dataset, along with the prompts listed in Tables 1 and 2, is used exclusively to assess the models’ prompt-following capabilities, as measured by the CLIP-T metric. These prompts include both stylistic and recontextualization examples, enabling evaluation across multiple dimensions of prompt adherence.
In contrast, the Unsplash partition and the prompts provided in Table 3 are employed to evaluate our proposed metric, R-FaceSim. These prompts are specifically crafted to encourage the generation of images in which the subject is directly facing the camera, with no facial occlusions, ensuring optimal conditions for facial similarity assessment.
@@ Appendix E Details on Evaluation Datasets and Prompts
Appendix F Details on Evaluation Metrics
CLIP-T.
CLIP-T measures how well the generated image aligns with the textual prompt. It is computed as the cosine similarity between the CLIP embeddings of the input prompt and the generated image. We use the CLIP ViT-B/32 model [22] to obtain the embeddings. A higher CLIP-T score indicates stronger prompt adherence.
Conventional FaceSim.
Conventional FaceSim metric is computed by extracting facial embeddings from the generated images using an Inception-ResNet [28] model pre-trained on VGGFace2 [2], followed by calculating the cosine similarity between these embeddings and those of the real subject. A discussion on the shortcomings of the conventional FaceSim metric, and the proposed improved version is given in the experiments section of the main paper.
R-FaceSim. Algorithm 3 outlines the computation of the Robust Face Similarity Metric (R-FaceSim), which leverages multiple images per identity, carefully curated prompts, and a generative text-to-image model. For each identity, one image is designated as the reference, while the remaining images constitute the test set. The model is either conditioned or fine-tuned on the reference image. For each prompt, the model generates an output image, from which the face is cropped and its identity embedding is extracted. This embedding is then compared—using cosine similarity—to the embeddings of faces from the corresponding test images. The mean similarity for a given prompt and identity is calculated by averaging these cosine similarities. The final R-FaceSim score is obtained by averaging these values across all identities and prompts. The models employed in this process are identical to those used in the Conventional FaceSim evaluation.
Input:
$I$ Set of identities with multiple images
$P$ Set of carefully chosen prompts
$M$ Generative text-to-image model
$C$ Face cropper
$E$ Identity embedding extractor
Output: $R\text{-}FaceSim$: Robust Face Similarity Score
for each $i\in I$ do
       $ref\leftarrow$ select_one($i$);
       $T\leftarrow i\setminus\{ref\}$;
      $M\leftarrow$ fine-tune/image-prompt($M$, $ref$);
      for each $p\in P$ do
             $g\leftarrow M(p)$;
             $f_{g}\leftarrow C(g)$;
             $e_{g}\leftarrow E(f_{g})$;
            $sumsim\leftarrow 0$, $count\leftarrow 0$;
             for each $t\in T$ do
                   $f_{t}\leftarrow C(t)$;
                   $e_{t}\leftarrow E(f_{t})$;
                   $sim\leftarrow cosine(e_{g},e_{t})$;
                   $sumsim\leftarrow sumsim+sim$;
                   $count\leftarrow count+1$;
            $S(i,p)\leftarrow\frac{sumsim}{count}$;
$R\text{-}FaceSim\leftarrow mean\{S(i,p):i\in I,p\in P\}$;
return $R\text{-}FaceSim$
Algorithm 3 Robust Face Similarity Metric (R-FaceSim)
Table 4: Comparison between the Robust FaceSim and FaceSim metrics, with percentage differences highlighted in cases of significant performance drops. All scores are averaged across both male and female identity classes.
Model Base Model Face Sim $\uparrow$ Robust Face Sim $\uparrow$ Relative Difference (%)
InstantID [29] SD-XL [21] 80.17 71.33 -11.0%
SD-XL Lightning [18] 82.91 74.77 -9.8%
PhotoMaker [17] SD-XL 67.28 67.16 -0.2%
SD-XL Lightning 66.67 63.02 -5.5%
PuLID [9] FLUX.1-dev [15] 84.76 75.72 -10.7%
Rank-1 LoRA [10] FLUX.1-dev 64.32 61.41 -4.2%
+ augmented inputs FLUX.1-dev 76.77 76.15 -0.8%
Meta-LoRA FLUX.1-dev 79.45 77.16 -2.9%
Appendix G Extra Metric Results
Table 5 presents a comparison of Meta-LoRA against state-of-the-art models from the literature, evaluated using the CLIP-T and R-FaceSim metrics. Meta-LoRA achieves the highest performance in facial similarity (R-FaceSim), establishing a new state-of-the-art in this aspect. For prompt adherence (CLIP-T), it ranks second, following PhotoMaker [17]. Notably, Meta-LoRA offers a more balanced trade-off between identity preservation and prompt alignment, whereas PhotoMaker tends to prioritize prompt fidelity at the expense of facial similarity.
Table 5: Quantitative comparison of our model with the state-of-the-art publicly available studies in the literature. The scores are taken for five seeds for both ’male’ and ’female’ subsets, and averaged to reach the final output. While CLIP-T is calculated from the Meta-PHD-FFHQ subset, R-FaceSim is computed using the Meta-PHD-Unsplash images. The 1st highest score is highlighted with blue, the 2nd with red, and the 3rd with green.
Model Base Model Train Dataset CLIP-T (%) $\uparrow$ R-FaceSim (%) $\uparrow$
InstantID [29] SD-XL [21] 60M 29.37 72.09
SD-XL Lightning [18] 30.52 75.26
PhotoMaker [17] SD-XL 112K 31.51 67.57
SD-XL Lightning 31.92 63.51
PuLID [9] FLUX.1-dev [15] 1.5M 30.95 75.72
Rank-1 LoRA FLUX.1-dev 0 31.63 61.41
+ augmented inputs 31.47 76.15
Meta-LoRA (ours) FLUX.1-dev 8K (male), 21K (female) 31.66 77.16
Appendix H Analysis on the Male Class Performance
The left panel of Figure 5 illustrates the distribution of R-FaceSim and CLIP-T scores for the ’male’ class across various hyperparameter configurations. Meta-LoRA consistently outperforms all baseline and previous state-of-the-art models, maintaining higher median and upper-bound scores. Compared to the female class, boundaries between different architectures (i.e., the performance gap) is easier to observe. Interestingly, even the baseline LoRA model demonstrates improved facial similarity compared to PuLID, suggesting that simpler adaptation methods can be competitive under certain conditions.
In Figure 5 (right), the plots show how CLIP-T and R-FaceSim scores vary with training iteration count. While the trends are similar to those observed for the female class (refer to the Quantitative Results section in the main paper), Meta-LoRA achieves a greater margin in R-FaceSim over prior state-of-the-art models when evaluated on the male subset.
Figure 5: Illustration of the metric score trends for Meta-LoRA and the standard LoRA [10] models on the Meta-PHD dataset, focusing on the ’male’ class. Left: A plot of R-FaceSim versus CLIP-T scores for all LoRA and Meta-LoRA models trained with ranks $\in$ 1, 2, 4, 8, 16 and across 250, 375, 500, 625, and 750 training iterations. The polynomial fit curves per each model are also shared for better comparison. Right: R-FaceSim and CLIP-T scores plotted against iteration count for rank-1 LoRA and Meta-LoRA models. These results are compared with state-of-the-art models from the literature, including PuLID [9], InstantID [29], and PhotoMaker [17]. To ensure a fair comparison, we use the FLUX.1-dev version of PuLID [15], which is consistent with the Meta-LoRA training setup. InstantID and PhotoMaker are based on SD-XL [21] and SD-XL Lightning [18], as public implementations of these models are not currently available for FLUX.1-dev.
Appendix I Performance Curves on Different Ranks
Figure 6 analyzes the evolution of CLIP-T and R-FaceSim scores across different fine-tuning iterations (i.e., 250, 375, 500, 625, 750) and selected rank values (i.e., 1, 4, and 16). For both gender classes, increasing the rank generally leads to improved facial similarity (higher R-FaceSim) but at the cost of reduced prompt adherence (lower CLIP-T), indicating a tendency toward overfitting at earlier stages of training.
While CLIP-T scores show a noticeable gap between different ranks across all iteration counts, the variation in R-FaceSim narrows as the number of iterations increases for both Meta-LoRA and LoRA+aug. These trends suggest that lower ranks are better suited for achieving a balanced trade-off between identity preservation and prompt following, particularly in the context of final model evaluations.
Figure 6: Visualization of CLIP-T and R-FaceSim performance for Meta-LoRA and baseline models across different fine-tuning iteration counts. The plots include results for both gender classes and three rank settings (i.e., 1, 4, and 16).
Appendix J Qualitative Results
Figures 7 and 8 demonstrate extended comparison for a diverse set of inputs and reference images.
Figure 7: Qualitative comparisons for an extended list of baselines for woman identities.
Figure 8: Qualitative comparisons for an extended list of baselines for man identities.